OpenStudy (anonymous):

Perform the indicated operation and simplify as much as possible: (2xy^4)^-2 divided by 3(x^4y)^-1

OpenStudy (anikay):

I'll get on this one in a sec

OpenStudy (anonymous):

thanks!!

OpenStudy (anikay):

oh wonderful, division of negative exponents, so really, we're dividing the second by the first....

OpenStudy (anonymous):

okay

OpenStudy (anikay):

\[\frac{ 3x^43y }{(2xy^4)^2} \]

OpenStudy (anikay):

\[\frac{3x^43y}{ 4x^2y^8 }\]

OpenStudy (anikay):

then common factors

OpenStudy (anikay):

\[\frac{ 3 }{ 4 } * \frac{ x^2 }{ y^4 }\] if I did my math right.

OpenStudy (anonymous):

wouldn't the 3 stay in the denominator\[3(x^4y)^{-1}={3 \over (x^4y)}\]

OpenStudy (anonymous):

so when you divide by this it is \[x^4y \over 3\]

OpenStudy (anikay):

you sure? maybe I read the problem wrong?

OpenStudy (anonymous):

x^2 divided by 12y^7 is the answer but idk how to get it

OpenStudy (anonymous):

Is it (3x^4y)^-1 or 3(x^4y)^-1

OpenStudy (anikay):

then chme is right

OpenStudy (anonymous):

\[(2xy^4)^{-2} \over 3(x^4y)^{-1}\] is this the question?

OpenStudy (anikay):

I did the work right, just messed up the 3

OpenStudy (anonymous):

\[\huge (x^a)^b=x^{ab}\]\[\huge \frac{ x^a }{ x^b }=x^{a-b}\]Things to memorize

OpenStudy (anikay):

^yes.... that's what I used right there

OpenStudy (anonymous):

chme yes that is the correct problem

OpenStudy (anonymous):

ChmE is right

OpenStudy (anikay):

yuppers, got that

OpenStudy (anonymous):

chme come back!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anikay):

(x^4*3y)/3(2x*y^4)^2

OpenStudy (anonymous):

\[(2xy^4)^{-2} \over 3(x^4y)^{-1}\]\[\frac{ x^4y }{ 3(2xy^4)^2 }\]\[\frac{ x^4y }{ 3(4x^2y^8) }\]\[\frac{ x^4y }{ 12x^2y^8 }\]\[\frac{ x^2 }{ 12y^7 }\]^Final Answer^

OpenStudy (anikay):

correct

OpenStudy (anikay):

medal

OpenStudy (anonymous):

yes thank you you are the best!!

OpenStudy (anonymous):

oh stop it